OpenStudy (anonymous):

My son is taking Physics and has this question and I have no idea how to help him since I don't understand it either. He needs to use the Coulomb's Law to figure it out is all I know. A thundercloud has an electric charge of 43.2 C near the top of the cloud and -38.7 C near the bottom of the cloud. The magnitude of the electric force between these two charges is 3.95x10^6 N. What is the average separation between these charges? (kc=8.99x10^9 N*m^2/C^2)

OpenStudy (irishboy123):

model it as if there were 2 point charges using Coulumb's inverse square law, F = k•q1•q2/d^2. F = force due to electro static charges k = Coulumb constant q1, q2 = the 2 point charges (remember the signs) you have F, k, q1, q2, so you just need to calc d, the distance between the 2 point charges.

OpenStudy (anonymous):

Did Irishboy cover it for you?

OpenStudy (anonymous):

sorry i had to step away just got back let me look it over a sec

OpenStudy (anonymous):

no i don't understand i know its something like this \[\sqrt{(8.99x10^9Nm ^2/c ^2) (4.32C)(-3.95C)}\div3.95x10^6N\] but im not sure how to do the calculations

OpenStudy (irishboy123):

F = (k x q1 x q2)/d^2 so d^2 = (k x q1 x q2)/F so d = sqrt { ( k x q1 x q2) / F }

OpenStudy (anonymous):

huh ok im lost isn't that what i have?

OpenStudy (anonymous):

how do I go from what I have to the answer is what I need, all the n, m and c throw me

OpenStudy (irishboy123):

\[d = \sqrt{\frac{ k * q1 * q2 }{ F }}\] this is \[\sqrt{\frac{ (8.99*10^{9} )* (43.2) * (-38.7) }{ -3.95*10^{6} }}\] if i have read the question correctly, giving 1950m as the answer....

OpenStudy (anonymous):

why is the denominator a negative?

OpenStudy (anonymous):

i think i kept forgetting to square it after the calculations thank you

OpenStudy (irishboy123):

force is negative because attractive force between +ve and -ve charges. but this is just a sign convention to make the equations work.